# Real Time Proportional Throughput Maximization: How much advance notice should you give your scheduler?

### v1

@@ 1.1 Abstract
We will be exploring a generalization of real time scheduling problem sometimes called the real time throughput maximization problem. Our input is a sequence of jobs specified by their release time, deadline and processing time. We assume that jobs are announced before or at their release time. At each time step, the algorithm must decide whether to schedule a job based on the information so far. The goal is to maximize the value of the sum of the processing times of jobs that finish before their deadline, or the total “active” time.
We extend this problem by defining a notion of $t$-advance-notice, a measure of how far in advance each job is given relative to their processing time. We show that there exists a $\frac{t}{2t+1}$-competitive algorithm when all jobs have $t$-advance-notice for $t\in[0,1]$. We also show that this ratio is optimal for all algorithms with $t$-advance-notice and that the upper bound of $\frac{t}{2t+1}$-competitiveness holds for all $t$, in particular that regardless of how much advance-notice is given, no algorithm can reach $\frac{1}{2}$-competitiveness.
@@ 1.2 Our model
Our input is a sequence of jobs $\mathcal{J}=(J_{1},\ldots,J_{n})$ where $J_{i}=(a_{i},r_{i},p_{i},d_{i})\in\mathbb{R}_{\geq 0}^{4}$. In this model, each job is announced by an adversary at time $a_{i}=a(J_{i})$, at this time, its “release time” $r_{i}=r(J_{i})$, processing time $p_{i}=p(J_{i})$ and deadline $d_{i}=d(J_{i})$ are revealed to the algorithm. The algorithm does not know what jobs may be announced in the future or the number of jobs $n$. At each time step, the algorithm ALG must decide whether to schedule a job based on the information so far, subject to the restriction that if a job $J_{i}$ is scheduled at time $s_{i}=s(J_{i})=s^{\textsc{ALG}}(J_{i})$ then no other job can be scheduled during the interval $s_{i},s_{i}+p_{i})$, furthermore, $r_{i}\leq s_{i}\leq s_{i}+p_{i}\leq d_{i}$. The goal is to maximize the value of the sum of the processing times $p_{i}$ of jobs that finish before their deadline, or the total “active” time. Later in this paper, we will also briefly explore the alternative goal of maximizing the total number of jobs that finish before their deadline (regardless of length), we call this case “unweighted”. Both these objectives are special cases of a “weight” maximization where we aim to maximize the total “weight” of jobs that finish before the deadline, in the former case, we assume that the weight is proportional to the processing time $p_{i}$ of each job, and in the latter we assume the weight is some fixed constant. We will also briefly look at maximizing the value of other weight functions such as $C$-benevolent (convex increasing) functions.
@@ 1.3 Related Works
The problem of real time throughput and it’s applications has been studied quite thoroughly.
@@
It is not difficult to show that in general, any algorithm can be made to perform arbitrarily bad. After all, an adversarial scheduler could schedule a job of weight 1, then schedule another job of arbitrarily large weight overlapping with that job if and only if the algorithm takes the original job. Regardless of what the algorithm does, its performance when compared to an optimal offline algorithm is arbitrarily large. For this reason, most of the existing work related to this problem tackle special cases or restrictions on the problem.
@@
The problem of job scheduling with fixed start and end times has been studied by Woeginger [WOEGINGER19945] where he found that maximizing the total weights of jobs could be done by a $\frac{1}{4}$-competitive algorithm when the weights are assumed to be given by functions he describes as $C$-benevolent and $D$-benevolent and the algorithm is allowed to “revoke” another job, stopping it’s execution early and forfeiting it’s value in exchange for scheduling another, better job. Notably, since the identity function (and any linear function) is $C$-benevolent, his result applies to the proportional case as well. The case where start and end times are fixed is often referred to as “interval scheduling”, as opposed to “job scheduling”. Later, still focusing on intervals Fung et al. [Fung2014] gives a $\frac{1}{2}$-competitive ratio barely random algorithm for “monotone” intervals as well as $C$-benevolent and $D$-benevolent weight functions. He also give a $\frac{1}{3}$-competitive algorithm for equal length jobs allowing for variable start times (also known as slack). In [Fung2014], he assumes that algorithms have the ability to “restart” other jobs, which is to say interrupt them similar to “revoking”, only now we also allow jobs that are revoked to be started again from scratch so long as their deadline has not past. This work built on previous results from Chrobak et al. [ChrobakRestarts] who showed that for equal processing time jobs deterministic algorithms a competitive ratio of $\frac{2}{3}$ is tight when restarting is allowed for unweighted job scheduling. However, they also show that when inputs are given sufficiently large “slack”, restarting is not needed for their algorithm to reach the desired competitive ratio. The more general case of restarting algorithms for unweighted job scheduling is studied by Hoogeveen et al. [HOOGEVEEN2000193] where he gives a tight $\frac{1}{2}$ competitive ratio for arbitrarily processing times in the unweighted restart model.
@@
On the offline side of the problem, [BaptisteOn4] Baptiste gives a polynomial time and space algorithm for maximizing the number of on time jobs. As for the weighted case, the very closely related task of minimizing weight of late jobs is shown to be NP-Complete by Karp [Karp1972]. A $(1-\frac{1}{e}-\varepsilon)$-approximation is given by Karakostas [KARAKOSTAS2025115258]. Meanwhile, Borodin et al. [Borodin2003] show that there exists a priority algorithm with $\frac{1}{3}$-competitive ratio for unweighted job scheduling and $\frac{1}{4}$-competitive ratio for proportional weights.
@@
In general, when studying real time scheduling, the time of announcement $a_{i}$ and the time of release $r_{i}$ are assumed to be the same for all jobs. That is to say, to the best of our knowledge, we are the first to explore a model where jobs may be announced ahead of their release time in a real-time model. The main benefit of this addition is that it allows us to reach a finite, and relatively good, competitive ratio without the need to restart or preempt. The problem of throughput in a real time context is very well motivated by many applications such as CPU scheduling,among other things. In many contexts, however, it may be inappropriate or detrimental to stop or revoke a job in the middle of it’s execution. Somebody who booked a room, or a service, may for example find it inconvenient to lose access in the middle of their alloted time. It may, in these circumstances, seem far more reasonable to require users or “jobs” to announce their availability and required time in advance. Ultimately, this extension feels very natural, and provides what we view as an appropriate alternative to the preemption models studied elsewhere.
@@ 1.4 Preliminary
It is clear that in general, the inclusion of an announcement time does not help us. Since giving advance notice $a_{i}<r_{i}$ does not impact the value of the optimal offline solution (which knows about every job from the very beginning), an adversary designing the sequence of jobs $\mathcal{J}$ can always make an algorithm perform worst by making $a_{i}=r_{i}$ without impacting the optimal solution. We therefore need to place a restriction on the value of $a_{i}$ to allow our algorithms to be able to reliably use it.
@@
2 The power and limitations of advance-notice
@@ 2.1 The power of advance-notice
In other words, at each announcement, run Off on the set of known jobs not run yet, possibly cutting off their release time to only the feasible times.
@@
We note that none of these jobs can be contained completely with $J_{i}$. This means that at most 2 jobs from any of these sets cover the left and right endpoints of $J_{i}$’s processing interval $s_{j},s_{j}+p_{j})$. Furthermore we note that $A_{i}$ can exist if no processes the left side, otherwise $A_{i}$ could simply be scheduled by $\mathcal{A}_{\textsc{Off}}$ to get a better schedule. As such since $t^{-1}\cdot p(J_{i})\geq p(J_{i})$. $p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})\leq t^{-1}\cdot p(J_{i})+p(J_{i})$.
@@ 2.2 The limitations of advance-notice
Suppose a sequence of jobs have $t$-advance notice, then no algorithm can achieve a better worst case asymptotic competitive ratio than $\frac{t}{2t+1}$ without preemption for proportional weight real time throughput maximization.
@@
If $J_{1}$ is scheduled at time $s_{1}$ then at $s_{1}$ we announce $J_{2}$ with $a_{2}=s_{1}+\gamma$, $r_{2}=s_{1}+1-2\cdot\gamma$, $p_{2}=\frac{p_{1}-\gamma}{t}$ and $d_{2}=r_{2}+p_{2}$. Note that $J_{2}$ is impossible to schedule since no preemptions is permitted. Next the adversary announces $J_{3}$ with $a_{3}=s_{1}+\gamma$, $r_{3}=s_{1}+2\cdot\gamma$ $p_{3}=\frac{\gamma}{t}$ $d_{3}=r_{3}+p_{3}$. The adversary will also announce: $J_{i}$ with $a_{i}=s_{1}+\gamma$, $r_{i}=d_{i-1}$, $p_{i}=d_{i}-r_{i}$ and $d_{i}=\min(r_{i}+\frac{\gamma}{t},r_{2})$. For $i>3$ until the interval $s_{1},s_{1}+p_{1}]$ is covered.
@@
So $\textsc{OPT}\cdot\frac{t}{2t+1}=\textsc{ALG}-\frac{(2\gamma+\gamma\cdot t)\cdot}{t\cdot(2t+1)}$. By our choose of $\gamma$ we get that $\textsc{OPT}\cdot\frac{t}{2t+1}+\varepsilon\leq\textsc{ALG}$
@@
The algorithm presented in theorem 2.1 is optimal on $t\in[0,1]$
Figure 3: ALGALG vs OPTOPT schedule
@@ 3 Beyond Proportionality:
The section above allows us to understand the competitive ratio of algorithms when we are given $t$-advance-notice when weights are proportional. While these results are quite strong, restricting ourselves to proportional weights might be in some situations be too strict. We will in this section show that the performance of algorithms for cases where we are not dealing with proportional weights and show that in some of the more obvious other cases (unweighted and $C$-benevolent) regardless of $t$, an algorithm running on an set of jobs with $t$-advance-notice performs arbitrarily poorly and cannot achieve a constant competitive ratio. We will then show that if we restrict announcement times such that $a_{i}<a_{j}\implies d_{i}<d_{j}$, we can achieve $\frac{1}{2}$ competitive ratio when jobs are unweighted.
@@ 3. C3)
In general, many of the nice properties of linear functions carry over to $C$-benevolent functions. In fact, many of the results from previous papers on interval selection and throughput apply to $C$-benevolent functions. It is why it is surprising that while we can get a finite competitive ratio with proportional weights, we cannot when the sequence of jobs is $f$-related to an arbitrary $C$-benevolent functions $f$.
@@
Any algorithm ALG can have arbitrarily bad competitive ratio when the a sequence of jobs is $f$-related to a $C$-benevolent function $f$ when jobs have $t$-advance-notice for $t\in(0,1]$.
@@
The proof for this is straightforwards, we once again use an adversarial argument. Our adversary chooses $f(x)=x^{\log_{\frac{1-\varepsilon}{t}}(N)}$ to be our function $f$ for some very small epsilon. We note that this function is $C$-benevolent for $N$ sufficiently larger than $t$.
@@ 3.2 Advance-notice for unweighted case
We call a sequence of jobs “unweighted” if it is $f$-related to the function $f(x)=1$.
@@
Suppose a sequence of jobs is unweighted and has $t$-advanced notice. Then no algorithm $ALG$ can achieve a finite competitive ratio without preemption
@@
We then announce $N$ jobs of arbitrarily small size such that they are contained within $J_{1}$’s interval. Since all jobs have the same weight, we can make the jobs processing time small enough that the start times of all of them, along with their end times and their announcement times are all contained within $r_{1},d_{1}]$.
@@ 4 Conclusion, Open Problems and Extensions
In the final section we showed that advance-notice does not help us much with some of the more “obvious” weight functions that people consider. These result seem to stem from a limitation of our definition of advance-notice. After all, by restricting our $(r_{i}-a_{i})$ to being proportional to $p_{i}$, it is clear that we gain information both about the processing time, but also about the weight of the job. As such, an interesting open question would be to see if we could get better results with a different definition of advance-notice that correspond more closely to the weight function used.
@@ 4.1 Acknowledgement
We are very appreciative of the feedback and encouragement from Professor Allan Borodin, and Changdao He who have provided me with valuable feedback and discussions throughout the construction of this paper.

### v2

@@ 1.1 Abstract
We will be exploring a generalization of real time scheduling problem sometimes called the real time throughput maximization problem. Our input is a sequence of jobs specified by their release time, deadline and processing time. We assume that jobs are announced before or at their release time. At each time step, the algorithm must decide whether to schedule a job based on the information so far. The goal is to maximize the value of the sum of the processing times of jobs that finish before their deadline, this is often called real time throughput with proportional weights.
We extend this problem by defining a notion of $t$-advance-notice, a measure of how far in advance each job is announced relative to their processing time. We show that there exists a $\frac{t}{2t+1}$-competitive algorithm when all jobs have $t$-advance-notice for $t\in[0,1]$, this gives us a competitive ratio of $\frac{1}{3}$ when $t$ is greater than or equal to $1$. We also show that this ratio is optimal for all algorithms with $t$-advance-notice and that the upper bound of $\frac{t}{2t+1}$-competitiveness holds for all $t$, in particular that regardless of how much advance-notice is given, no algorithm can reach $\frac{1}{2}$-competitiveness.
@@ 1.2 Our model
Our input is a sequence of jobs $\mathcal{J}=(J_{1},\ldots,J_{n})$ where $J_{i}=(a_{i},r_{i},p_{i},d_{i})\in\mathbb{R}_{\geq 0}^{4}$. In this model, each job is announced by an adversary at time $a_{i}=a(J_{i})$, at this time, its “release time” $r_{i}=r(J_{i})$, processing time $p_{i}=p(J_{i})$ and deadline $d_{i}=d(J_{i})$ are revealed to the algorithm. The algorithm does not know what jobs may be announced in the future or the number of jobs $n$. At each time step, the algorithm Alg must decide whether to schedule a job based on the information so far, subject to the restriction that if a job $J_{i}$ is scheduled at time $s_{i}=s(J_{i})=s^{\textsc{Alg}}(J_{i})$ then no other job can be scheduled during the interval $s_{i},s_{i}+p_{i})$, furthermore, $r_{i}\leq s_{i}\leq s_{i}+p_{i}\leq d_{i}$. The goal is to maximize the value of the sum of the processing times $p_{i}$ of jobs that finish before their deadline, or equivalently the total time any job is running. Later in this paper, we will also briefly explore the alternative goal of maximizing the total value of the “weights” of jobs that finish before their deadline. One examples is if we assign each weight to be a constant, we call this case “unweighted”. We will also briefly look at maximizing the value of other weight functions such as $C$-benevolent (convex increasing) functions.
@@ 1.3 Related Works
The problem of real time weighted throughput and it’s applications has been studied quite thoroughly.
@@
It is not difficult to show that in a real time online model, any algorithm can be made to perform arbitrarily bad. After all, an adversarial scheduler could schedule a job of weight 1, then schedule another job of arbitrarily large weight overlapping with that job if and only if the algorithm takes the original job. Regardless of what the algorithm does, its performance when compared to an optimal offline algorithm is arbitrarily large. For this reason, most of the existing work related to this problem tackle special cases or restrictions on the problem.
@@
The problem of job scheduling with fixed start and end times, called “interval” scheduling, has been studied by Woeginger [WOEGINGER19945] where he found that maximizing the total weights of jobs could be done by a $\frac{1}{4}$-competitive algorithm when the weights are assumed to be given by functions he describes as $C$-benevolent and the algorithm is allowed to “revoke” another job, stopping it’s execution early and forfeiting it’s value in exchange for scheduling another, better job. Notably, his result applies to the proportional case as well. Later, still focusing on intervals Fung et al. [Fung2014] gives a $\frac{1}{2}$-competitive ratio barely random algorithm for “monotone” intervals as well as $C$-benevolent and $D$-benevolent weight functions. He also give a $\frac{1}{3}$-competitive algorithm for equal length jobs allowing for variable start times (also known as slack). In [Fung2014], he assumes that algorithms have the ability to “restart” other jobs, which is to say interrupt them similar to “revoking”, only now we also allow jobs that are revoked to be started again from scratch so long as their deadline has not past. Chrobak et al. [ChrobakRestarts] also studied this model and showed that for equal processing time jobs deterministic algorithms a competitive ratio of $\frac{2}{3}$ is tight when restarting is allowed for unweighted job scheduling. However, they also show that when inputs are given sufficiently large “slack”, restarting is not needed for their algorithm to reach the desired competitive ratio. The more general case of restarting algorithms for unweighted job scheduling is studied by Hoogeveen et al. [HOOGEVEEN2000193] where he gives a tight $\frac{1}{2}$ competitive ratio for arbitrarily processing times in the unweighted restart model. Changdao He[he2025revokevsrestartunweighted] shows that in contrast to restarting, revoking on it’s own is not enough to get a constant competitive ratio for job scheduling with unweighted jobs.
@@
On the offline side of the problem, [BaptisteOn4] Baptiste gives a polynomial time and space algorithm for maximizing the number of on time jobs when jobs are allowed to stop and resume. As for the case without resuming, the very closely related task of minimizing weight of late jobs is shown to be NP-Complete by Karp [Karp1972]. A $(1-\frac{1}{e}-\varepsilon)$-approximation is given by Karakostas [KARAKOSTAS2025115258]. Meanwhile, Borodin et al. [Borodin2003] show that there exists a priority algorithm with $\frac{1}{3}$-approximation ratio for unweighted job scheduling and $\frac{1}{4}$-approximation ratio for proportional weights, they also show that no fixed-priority algorithm can achieve better approximation ratio than $\frac{1}{2}$ and $\frac{1}{4}$ for the unweighted and proportional cases respectively.
@@
In general, when studying real time scheduling, the time of announcement $a_{i}$ and the time of release $r_{i}$ are assumed to be the same for all jobs. That is to say, to the best of our knowledge, we are the first to explore a model where jobs may be announced ahead of their release time in a real-time model. The main benefit of this addition is that it allows us to reach a finite, competitive ratio without the need to restart or revoke. The problem of throughput in a real time context is very well motivated by many applications such as CPU scheduling,among other things. In many contexts, however, it may be inappropriate or detrimental to stop or revoke a job in the middle of it’s execution. Somebody who booked a room, or a service, may for example find it inconvenient to lose access in the middle of their alloted time. It may, in these circumstances, seem far more reasonable to require users or “jobs” to announce their availability and required time in advance. Ultimately, this extension feels very natural, and provides what we view as an appropriate alternative to the preemption models studied elsewhere.
@@ 1.4 Preliminary
In general, the inclusion of an announcement time does not necessarily help us against an adversary relative to a model where jobs are announced at their release times. Since giving advance notice $a_{i}<r_{i}$ does not impact the value of the optimal offline solution (which knows about every job from the very beginning), an adversary designing the sequence of jobs $\mathcal{J}$ can always make an algorithm perform worst by making $a_{i}=r_{i}$ without impacting the optimal solution. We therefore need to place a restriction on the value of $a_{i}$ to allow our algorithms to be able to reliably use it.
@@
2 The power and limitations of advance-notice for proportional weights
@@ 2.1 The power of advance-notice
In other words, at each announcement, run Off on the set of known jobs not run yet, possibly increasing their release time to only the feasible times.
@@
We note that none of these jobs can be contained completely with $J_{i}$. This means that at most 2 jobs from any of these sets cover the left and right endpoints of $J_{i}$’s processing interval $s_{j},s_{j}+p_{j})$. Furthermore we note that $A_{i}$ can exist if no processes the left side, otherwise $A_{i}$ could simply be scheduled by $\mathcal{A}_{\textsc{Off}}$ to get a better schedule. As such since $t^{-1}\cdot p(J_{i})\geq p(J_{i})$ when $t\in[0,1]$. $p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})\leq t^{-1}\cdot p(J_{i})+p(J_{i})$.
@@
Since by definition, if a sequence of jobs has $t$-advance-notice, it has $t^{\prime}$-advance-notice for $t^{\prime}<t$ we get the following corollary:
Corollary 2.1.
There exists an online algorithm Alg which is at least $\min\{\frac{t}{2t+1},\frac{1}{3}\}$-competitive for real time proportional throughput when all jobs have $t$-advance-notice for any $t$
@@ 2.2 The limitations of advance-notice
Suppose a sequence of jobs have $t$-advance notice, then no algorithm can achieve a better worst case asymptotic competitive ratio than $\frac{t}{2t+1}$ without the ability to revoke previously scheduled jobs for proportional weight real time throughput maximization.
@@
If $J_{1}$ is scheduled at time $s_{1}$ then at $s_{1}+\gamma$ we announce $J_{2}$ with $a_{2}=s_{1}+\gamma$, $r_{2}=s_{1}+1-2\cdot\gamma$, $p_{2}=\frac{p_{1}-\gamma}{t}$ and $d_{2}=r_{2}+p_{2}$. Note that $J_{2}$ is impossible to schedule since no preemptions is permitted. Next the adversary announces $J_{3}$ with $a_{3}=s_{1}+\gamma$, $r_{3}=s_{1}+2\cdot\gamma$ $p_{3}=\frac{\gamma}{t}$ $d_{3}=r_{3}+p_{3}$. The adversary will also announce: $J_{i}$ with $a_{i}=s_{1}+\gamma$, $r_{i}=d_{i-1}$, $p_{i}=d_{i}-r_{i}$ and $d_{i}=\min(r_{i}+\frac{\gamma}{t},r_{2})$. For $i>3$ until the interval $s_{1},s_{1}+p_{1}]$ is covered.
@@
So $\textsc{Opt}\cdot\frac{t}{2t+1}=\textsc{Alg}-\frac{(2\gamma+\gamma\cdot t)\cdot}{t\cdot(2t+1)}$. By our choice of $\gamma$ we get that $\textsc{Opt}\cdot\frac{t}{2t+1}+\varepsilon\leq\textsc{Alg}$
Figure 3: ALGALG vs OPTOPT schedule
@@
The algorithm presented in theorem 2.1 has optimal competitive ratio on sequences of jobs with $t$-advance-notice and $t\in[0,1]$
@@ 3 Beyond Proportionality:
The section above allows us to understand the competitive ratio of algorithms when we are given $t$-advance-notice when weights are proportional. Despite these results, restricting ourselves to proportional weights might in some situations seem too strict. We will in this section look at the performance of algorithms in cases where we are not dealing with proportional weights and show that in some of the more obvious other cases (unweighted and $C$-benevolent) regardless of $t$, an algorithm running on a set of jobs with $t$-advance-notice can be made to perform arbitrarily poorly and cannot achieve a constant competitive ratio.
@@ 3. C3)
In general, many of the nice properties of linear functions carry over to $C$-benevolent functions. It is why it may be surprising that while we can get a finite competitive ratio with proportional weights, we do not get that same guarantee if the sequence of jobs is $f$-related to an arbitrary $C$-benevolent functions $f$.
@@
For any $t\in(0,1)$, there exists a $C$-benevolent function $f$ and a sequence of $f$-related jobs such that any algorithm Alg can have arbitrarily bad competitive ratio against this sequence of jobs.
@@
We once again use an adversarial argument. Our adversary chooses $f(x)=x^{\log_{\frac{1-\varepsilon}{t}}(N)}$ to be our function $f$ for some very small epsilon. We note that this function is $C$-benevolent for $N$ sufficiently larger than $t$.
@@ 3.2 Advance-notice for unweighted case
We call a sequence of jobs “unweighted” if it is $f$-related with the function $f(x)=1$.
@@
For any $t$, there is a sequence of unweighted jobs that has $t$-advanced notice such that no algorithm $ALG$ can achieve a finite competitive ratio without preemption
@@
We then announce jobs $J_{2},\ldots,J_{N+1}$ such that $\forall i,j\in\{2,\ldots,N+1\}$: $a_{i},d_{i}]\subset(r_{1},d_{1})$ and $a_{i},d_{i}]\cap[a_{j},d_{j}]=\emptyset$. Since all jobs have the same weight, the value of a algorithm that schedules $J_{2},\ldots,J_{N+1}$ is $N$, as such Opt has value at least $N$ but $\textsc{Alg}=1$ as desired.
@@ 4 Conclusion, Open Problems and Extensions
In the final section we showed that advance-notice does not help us much with some of the more “obvious” weight functions that people consider. As such, an interesting open question would be to see if we could get better results with a different definition of advance-notice that correspond more closely to the weight function used.
@@ 4.1 Acknowledgement
We are very appreciative of the feedback and encouragement from Allan Borodin, and Changdao He who have provided valuable feedback and discussions throughout the construction of this paper.